# Game-Theoretic Regularized Self-Play Alignment of Large Language Models

### v1

@@ Abstract
Self-play alignment algorithms have been developed as effective methods for fine-tuning large language models (LLMs), formulating preference optimization as a two-player game. However, the regularization with respect to the reference policy, which is crucial for mitigating over-optimization, has been insufficiently investigated in self-play alignment. In this paper, we show that our regularization method can improve the unregularized self-play significantly. To study the impact of different regularizations in self-play alignment, we propose Regularized Self-Play Policy Optimization (RSPO). This generalized framework regularizes the self-play by simply adding a chosen regularization term into the loss while maintaining provable last-iterate convergence to the Nash Equilibrium of the corresponding regularized game. Surprisingly, empirical evaluations using the Mistral-7B-Instruct base model reveal that forward KL divergence regularization reduces response length in RSPO, whereas reverse KL divergence markedly improves raw win rates. RSPO with a linear combination of forward and reverse KL divergence regularization substantially increases the length-controlled win rate in AlpacaEval-2, elevating the unregularized self-play alignment method (SPPO) from $28.53\%$ to $35.44\%$. Finally, we show that RSPO also improves the response diversity.
@@ 1 Introduction
Figure 1: Our Regularized Self-Play Policy Optimization (RSPO) with Mistral-7B-Instruct outperforms Llama-3-70B, SimPO (Meng et al., 2024), and SPPO (Wu et al., 2024) in AlpacaEval-2 with length-controlled (LC) win rate.
Large Language Models (LLMs) recently have obtained remarkable capabilities to accomplish a range of tasks (Jiang et al., 2023a; Dubey et al., 2024; DeepSeek-AI et al., 2025), generating more desirable and helpful content following the user’s intention. One of the most important methods to align LLMs with human intentions is Reinforcement Learning from Human Feedback (RLHF), maximizing a preference-based reward penalized by a reverse KL regularization term of LLM policy and a supervised fine-tuning (SFT) reference model (Christiano et al., 2017; Ouyang et al., 2022; Rafailov et al., 2024; Azar et al., 2024; Xiong et al., 2024). This regularization is crucial in RLHF to prevent over-optimization, which has been extensively studied and even extended beyond KL divergence (Go et al., 2023; Huang et al., 2024).
@@
Self-play is a general line of works conducting iterative self-competition of models, which has been demonstrated as an effective approach for improving AI systems (Goodfellow et al., 2020; Wang et al., 2022), particularly in strategic decision-making problems (Silver et al., 2016; Heinrich and Silver, 2016; Pinto et al., 2017; Brown and Sandholm, 2018). In the human alignment of LLMs, self-play recently started to be used and has shown superior empirical performance than other iterative RLHF methods on benchmarks like AlpacaEval and Arena-Hard Evaluation (Dubois et al., 2024; Jiang et al., 2024; Wu et al., 2024; Rosset et al., 2024). By formulating the preference optimization problem as a two-player game, self-play alignment methods seek to identify a Nash Equilibrium (NE) of the game in which utility is determined by a general preference model (Munos et al., 2023; Calandriello et al., 2024; Azar et al., 2024). This NE is regarded as the most aligned LLM policy, achieved without Bradley-Terry (BT) assumption (David, 1963).
@@
Despite the significant empirical improvements achieved through self-play, the impact of regularization to the reference policy—commonly used in RLHF to mitigate over-optimization—has received insufficient investigation in self-play alignment. Most existing self-play methods lack explicit regularization (Swamy et al., 2024; Rosset et al., 2024; Wu et al., 2024; Wang et al., 2024; Gao et al., 2024). In practice, unregularized self-play is also susceptible to over-optimization, particularly when the preference model is misspecified. While some approaches incorporate regularization, they are typically constrained to a reverse KL divergence penalty that restricts deviations from the reference policy (Munos et al., 2023; Zhang et al., 2024).
In this paper, we introduce a generalized framework for incorporating diverse regularization methods into self-play alignment, termed Regularized Self-Play Policy Optimization (RSPO). RSPO offers a simple way to apply various regularization strategies in self-play by directly adding the regularization term to the loss function, while maintaining last-iterate convergence to the NE of the corresponding regularized preference optimization game. Empirical analysis reveals distinct effects of different regularization methods: forward KL regularization reduces the response length in RSPO, whereas reverse KL regularization significantly enhances the raw win rate. Consequently, we adopt a linear combination of forward and reverse KL divergences, yielding a substantial improvement over the unregularized self-play alignment method, SPPO (Wu et al., 2024), on various benchmarks. Particularly on AlpacaEval-2, RSPO outperforms SPPO with a $6.9$ percentage points increase in length-controlled win rate (LCWR) and an $18$ percentage point LCWR improvement over the base model, Mistral-7B-Instruct. Furthermore, we offer an analysis of response diversity that regularization also promotes greater diversity. In summary, regularization plays a crucial role in self-play alignment, significantly improving both the quality and diversity of responses in previously unregularized self-play methods.
2 Related Work
Azar et al. (2024) introduced the first approach for optimizing general preference models. Nash-MD (Munos et al., 2023) pioneered the application of self-play to general preference optimization by framing it as a two-player game. Subsequent methods have either focused on learning the NE of the original unregularized game (e.g. (Swamy et al., 2024; Wu et al., 2024; Rosset et al., 2024; Wang et al., 2024)) or the NE of a reverse-KL-regularized preference optimization game (e.g. (Munos et al., 2023; Calandriello et al., 2024; Zhang et al., 2024)). In contrast, our work explores a broader class of divergence-based regularization techniques for self-play alignment.
We emphasize the distinction between our self-play approach and self-play methods based on pairwise comparisons, which construct loss functions by leveraging the difference in policy logits between preferred and rejected responses—such as Direct Policy Optimization (DPO) (Rafailov et al., 2024) and Identity Policy Optimization (IPO) (Calandriello et al., 2024). Direct Nash Optimization (DNO) (Rosset et al., 2024) and Iterative Nash Policy Optimization (INPO) (Zhang et al., 2024) follow the Mirror Descent (MD) update (Beck and Teboulle, 2003) while computing loss using pairwise comparisons. However, optimizing such pairwise-comparison-based losses can lead to only an increase in the relative likelihood gap without necessarily enhancing the absolute probability of the preferred response (Pal et al., 2024). In contrast, our method directly approximates the MD update by converting MD as a reinforcement learning problem, thereby circumventing the limitations of pairwise comparison-based approaches.
Online iterative RLHF, which incorporates a reliable reward or preference model—including self-play—functions as a self-improving framework by iteratively generating new data using models and optimizing policies based on this data (Schulman et al., 2017; Ouyang et al., 2022; Bai et al., 2022; Touvron et al., 2023; Dong et al., 2024). Moreover, extending powerful offline methods such as Direct Preference Optimization (DPO) to iterative frameworks has led to significant performance gains (Xu et al., 2023; Liu et al., 2023; Tran et al., 2023; Dong et al., 2024; Calandriello et al., 2024; Pang et al., 2024; Xiong et al., 2024; Guo et al., 2024; Tajwar et al., 2024; Cen et al., 2024; Xie et al., 2024). In contrast, our work investigates general preference optimization through self-play from a game-theoretic perspective, shifting the objective from conventional RL optimization to the computation of NE.
@@ 3 Preliminaries
We denote a prompt as $x$, a response as $y$, and a LLM policy as $\pi(y|x)$, where $\pi(\cdot|x)\in\Delta_{\mathcal{Y}}$, $\mathcal{X}$ is the set of all prompts and $\mathcal{Y}=\{y^{0},y^{1},\cdots\}$ is the set of all responses. We denote the probability simplex over the responses given a specific prompt as $\Delta_{\mathcal{Y}}$. We parametrize the LLM policy $\pi$ as $\pi_{\theta}$. The reference policy is an LLM denoted as $\mu\in\Delta^{\mathcal{X}}_{\mathcal{Y}}$. For notational brevity, we remove the dependence of policy $\pi$ and loss functions on the prompt $x$ throughout the paper.
@@ 3.1 Game-Theoretic Preference Optimization
We study the preference optimization problem in an online setting by formulating it as a two-player max-min game, as studied in previous self-play works (Wu et al., 2024). The players are two LLMs whose strategies are LLM policies, denoted as max-player $\pi$ and min-player $\pi^{\prime}$. The utility of the max-player is the preference:
@@
where $u:\Delta^{\mathcal{X}}_{\mathcal{Y}}\times\Delta^{\mathcal{X}}_{\mathcal{Y}}\rightarrow\mathbb{R}$ is linear in $\pi$ and $\pi^{\prime}$; $\mathbb{P}:\mathcal{X}\times\mathcal{Y}\times\mathcal{Y}\rightarrow[0,1]$ is a general preference model that quantifies the preference of $y$ over $y^{\prime}$ given a prompt. We extend the notation $\mathbb{P}(y\succ\pi^{\prime})=\mathbb{E}_{y^{\prime}\sim\pi^{\prime}}[\mathbb{P}(y\succ y^{\prime})]$ for any response $y$. The objective is finding a NE policy $\pi^{*}$ of the preference model:
@@
Therefore, an NE strategy $\pi^{*}$ is an LLM that can generate the most preferred responses in expectation, thus achieving human alignment based on the preference model.
Existing game-theoretic self-play methods solve this NE following Algorithm 1 (Wu et al., 2024; Swamy et al., 2024; Zhang et al., 2024; Wang et al., 2024). Specifically, the policy is first initialized as $\pi_{0}=\mu$. Then in each iteration $t$, the opponent is set to be the last-iterate policy $\pi_{t}$ (the reason why it is called self-play), and the responses are sampled from $\pi_{t}$ (Line 4). The pairwise preferences of the sampled responses are collected using the preference model $\mathbb{P}$ (Line 5). The policy parameters are updated by minimizing a specified loss function $\mathcal{L}(\theta;\mathbb{P})$ based on preferences over responses (Line 6). The loss function $\mathcal{L}(\theta;\mathbb{P})$ is dependent on the inherent online learning method.
@@ 3.2 Preference Optimization via Multiplicative Weights Update
An effective self-play method to solve the preference optimization game in Equation 2 is Self-Play Policy Optimization (SPPO) (Wu et al., 2024). SPPO derives its loss function from the iterative no-regret learning algorithm, Multiplicative Weights Update (MWU) (Freund and Schapire, 1997). Specifically in a game setting, denote learning rate as $\eta$, and normalization constant $Z(\pi_{t})$. In any iteration $t$, the policy update $\forall y\in\mathcal{Y}$ is
$\pi_{t+1}(y)=\pi_{t}(y)\cdot\frac{\exp\Big{(}\eta\mathbb{E}_{y^{\prime}\sim\pi_{t}}[u({y};y^{\prime})]\Big{)}}{Z(\pi_{t})},$ (3)
where $u(y;y^{\prime})$ is the utility function defined in Equation (1), with $y$ treated as a pure strategy.
The practical loss function of SPPO for policy update in each iteration $t$ is the square error between LHS and RHS in Equation 3 at a logarithmic scale, defined as
@@
SPPO converges to the NE of the preference optimization game in Equation 2. However, after running multiple iterations, the deviation of the policy $\pi_{\theta}$ from $\mu$ can be large. Such deviation is particularly problematic when the preference model is only accurate at evaluating responses sampled from the reference policy (Munos et al., 2023). Furthermore, in aligning LLMs in practice, the preference model is typically a surrogate $\hat{\mathbb{P}}$, such as PairRM (Jiang et al., 2023b), which may be misspecified at some out-of-distribution responses and inaccurate due to estimation error or limited model expressiveness (e.g., PairRM is only a 0.4B model), causing over-optimization problem. Regularizing the policy optimization to a reference SFT model, which is typically trained on high-quality data (Ouyang et al., 2022), can mitigate the problem. We provide a synthetic example in Appendix C.1 to demonstrate the problem.
@@ 3.3 Regularized Preference Optimization Game with Reference Policy
To address the regularization in self-play, we adopt the objective in Nash Learning from Human Feedback (Munos et al., 2023), and extend the KL divergence regularization to a general regularization function, to penalize the deviation from reference policy. We define a convex regularization function $R:\Delta^{\mathcal{X}}_{\mathcal{Y}}\times\Delta^{\mathcal{X}}_{\mathcal{Y}}\rightarrow(-\infty,\infty)$, where $R(\pi,\mu)$ measures the distance between $\pi$ and the reference model $\mu$, such as KL divergence $D_{\text{KL}}(\pi\|\mu)$. Denote regularization temperature as $\tau$, the objective becomes to optimize a regularized preference model by solving the NE $(\pi^{*},\pi^{*})$ of the regularized game, where the utility of max player is still $u(\pi;\pi^{\prime})=\mathbb{P}(\pi\succ\pi^{\prime})$:
@@
We provide proof of the existence and uniqueness of this NE in Appendix A.2. Various methods leverage Mirror Descent (MD) to find a regularized NE in Equation 5 (Munos et al., 2023; Calandriello et al., 2024; Zhang et al., 2024; Wang et al., 2024), based on its last-iterate convergence.
However, these MD-based methods are only compatible with the reverse KL divergence regularizer. Nash-MD [^1] Throughout the paper, regularization specifically refers to the one between $\pi$ and $\mu$, rather than between $\pi$ and $\pi_{t}$. addresses the reverse KL regularization of $\pi$ and $\mu$ using a geometric mixture policy $\pi_{t}^{\mu}$ defined as $\pi_{t}^{\mu}(y)={\pi_{t}(y)^{1-\eta\tau}\mu(y)^{\eta\tau}}/{\sum_{y^{\prime}}\pi_{t}(y^{\prime})^{1-\eta\tau}\mu(y^{\prime})^{\eta\tau}}$ (Munos et al., 2023), which updates policy as:
$\pi_{t+1}=\arg\min_{\pi}-\eta\langle\pi,\partial_{\pi}u(\pi_{t};{\pi^{\mu}_{t}})\rangle+D_{\text{KL}}(\pi,\pi_{t}^{\mu}).$ (6)
While the LLMs optimized via self-play exhibit significant improvement (Wu et al., 2024; Wang et al., 2024; Zhang et al., 2024), they all have limited regularization of $\pi$ and $\mu$. They either completely lack explicit regularization or only employ reverse KL divergence, imposing only a narrow form of regularization. The potential benefits of alternative regularization, such as adopting other $f$-divergences than reverse KL, remain unexplored.

### v2

# RSPO: Regularized Self-Play Alignment of Large Language Models

@@ Abstract
Self-play alignment has emerged as an effective approach for fine-tuning large language models (LLMs), formulating preference optimization as a two-player game. However, the regularization with respect to the reference policy, which is crucial for mitigating over-optimization, has been insufficiently investigated in self-play alignment. To study the impact of different regularization strategies, we propose Regularized Self-Play Policy Optimization (RSPO), a general and modular framework that unifies prior methods and enables simple plug-and-play integration of various regularizers, meanwhile preserving convergence to Nash equilibrium of the corresponding regularized game.Our empirical study involving over $120$ fine-tuned Mistral-7B-Instruct models reveals that forward KL divergence regularization reduces response length, whereas reverse KL divergence markedly improves raw win rates. Crucially, RSPO regularized with a linear combination of forward and reverse KL divergence significantly boosts the length-controlled win rate on AlpacaEval-2 from $28.5\%$ (unregularized self-play, SPPO) to $35.4\%$, and consistently demonstrates superior performance on Arena-Hard, MT-Bench, ArmoRM scores, and response diversity. Combining simplicity, convergence guarantees, and significant empirical gains, RSPO offers a strong foundation for exploring regularized self-play in language model alignment.
@@ 1 Introduction
Self-play is a line of work conducting iterative self-competition of models, which has been demonstrated as an effective approach for improving AI systems (Goodfellow et al., 2020; Wang et al., 2022), particularly in strategic decision-making problems (Silver et al., 2016; Heinrich and Silver, 2016; Pinto et al., 2017; Brown and Sandholm, 2018). In the human alignment of LLMs, self-play recently started to be used and has shown superior empirical performance than other iterative Reinforcement Learning from Human Feedback (RLHF) methods on popular benchmarks (Dubois et al., 2024; Jiang et al., 2024; Wu et al., 2024; Rosset et al., 2024). By formulating the preference optimization problem as a two-player game, self-play alignment methods seek to identify a Nash Equilibrium (NE) of the game in which utility is determined by a general preference model (Azar et al., 2024; Munos et al., 2023; Calandriello et al., 2024). This NE is regarded as the most aligned LLM policy achieved without Bradley-Terry (BT) reward modeling (David, 1963), which has shown under-performance compared to general preference modeling (Ye et al., 2024).
Figure 1: RSPO is flexible for general regularization. The estimation of Nash-MD policy update direction $\partial_{\pi}\mathbb{P}(\pi\succ\pi_{t}^{\mu})$ requires samples from geometric mixture policy $\pi_{t}^{\mu}$. Such update approach is only compatible with reverse KL divergence for regularization.
@@
Despite the significant empirical improvements achieved through self-play, the impact of regularization to the reference policy—commonly used in RLHF to mitigate over-optimization—has received insufficient investigation in self-play alignment. Most existing self-play methods completely lack explicit regularization (Wu et al., 2024; Rosset et al., 2024; Swamy et al., 2024; Wang et al., 2024a; Gao et al., 2024). In practice, unregularized self-play is also susceptible to over-optimization, particularly when the preference model is inaccurate or misspecified. Although a few recent self-play approaches like Nash-MD (Munos et al., 2023) incorporate reverse KL divergence as a regularization penalty (Calandriello et al., 2024; Wang et al., 2024a; Zhang et al., 2024), it remains unclear whether reverse KL is optimal for alignment, and the broader impact of alternative regularization strategies in self-play remains insufficiently explored. Moreover, the extension of current approaches to general forms of regularization is challenging, as their training protocols are intrinsically reliant on the reverse KL divergence for regularization Munos et al. (2023) (see Figure 1).
In this work, we introduce a flexible framework for incorporating diverse regularization methods into self-play alignment, termed Regularized Self-Play Policy Optimization (RSPO):
  • RSPO offers a simple way to apply various regularization strategies in self-play by directly adding the regularization term to our proposed unified self-play loss function, while maintaining last-iterate convergence to NE of the corresponding regularized preference optimization game.
  • We conduct empirical analysis on over $120$ fine-tuned Mistral-7B-Instrct models, which reveals distinct effects of different regularizations: forward KL regularization reduces the response length, whereas reverse KL regularization significantly improves the raw win rate.
  • Consequently, we propose a linear combination of forward and reverse KL divergence, yielding a substantial improvement over the unregularized self-play alignment method (SPPO (Wu et al., 2024)) on popular benchmarks: AlpacaEval-2.0, Arena-Hard-v0.1, and MT-bench. Particularly on AlpacaEval-2.0, RSPO outperforms SPPO with improvement of $6.9\%$ in length-controlled win rate (LCWR).
  • Finally, we provide quantitative evaluations on other aspects of the generated responses, including diversity via self-BLEU (Zhu et al., 2018), and instruction following, truthfulness, honesty, helpfulness via ArmoRM (Wang et al., 2024b). Regularization demonstrates improvements on almost all aspects of the responses, indicating the comprehensive effectiveness of regularization in self-play alignment.
@@ 2 Preliminaries
We denote a prompt as $x$, a response as $y$, and a LLM policy as $\pi(y|x)$, where $\pi(\cdot|x)\in\Delta_{\mathcal{Y}}$. We denote the set of all prompts as $\mathcal{X}$, and the set of all responses as $\mathcal{Y}=\{y^{0},y^{1},\cdots\}$. We use $\Delta_{\mathcal{Y}}$ to denote the probability simplex over the responses given a specific prompt. We parametrize the LLM policy $\pi$ as $\pi_{\theta}$. The reference policy is an LLM denoted as $\mu\in\Delta^{\mathcal{X}}_{\mathcal{Y}}$. For notational brevity, we remove the dependence of policy $\pi$ and loss functions on the prompt $x$ throughout the paper.
@@ 2.1 Game-Theoretic Preference Optimization
We study the preference optimization problem in an online setting by formulating it as a two-player max-min game, as studied in previous self-play works (Wu et al., 2024). The players are two LLMs whose strategies are LLM policies, denoted as max-player $\pi$ and min-player $\pi^{\prime}$. The utility of the max-player is expressed as the preference of itself over the min-player:
@@
where $u:\Delta^{\mathcal{X}}_{\mathcal{Y}}\times\Delta^{\mathcal{X}}_{\mathcal{Y}}\rightarrow\mathbb{R}$ is linear in $\pi$ and $\pi^{\prime}$; $\mathbb{P}:\mathcal{X}\times\mathcal{Y}\times\mathcal{Y}\rightarrow[0,1]$ is a general preference model that quantifies the preference of $y$ over $y^{\prime}$ given a prompt. We extend the notation $\mathbb{P}(y\succ\pi^{\prime})=\mathbb{E}_{y^{\prime}\sim\pi^{\prime}}[\mathbb{P}(y\succ y^{\prime})]$. The objective is finding a NE policy $\pi^{*}$ of the preference model:
@@
Therefore, an NE strategy $\pi^{*}$ is an LLM that can generate the most preferred responses in expectation, thus achieving human alignment based on the preference model. Most existing self-play alignment methods aim to solve this NE following Algorithm 1 (Wu et al., 2024; Rosset et al., 2024; Swamy et al., 2024; Wang et al., 2024a).
@@ 2.2 Preference Optimization via Multiplicative Weights Update
An effective self-play method to solve the preference optimization game in Equation 2 is Self-Play Policy Optimization (SPPO) (Wu et al., 2024). SPPO derives its loss function from the iterative no-regret learning algorithm, Multiplicative Weights Update (MWU) (Freund and Schapire, 1997). Specifically in a game setting, denote learning rate as $\eta$, and normalization constant $Z(\pi_{t})$. In any iteration $t$, the policy update $\forall y\in\mathcal{Y}$ is $\pi_{t+1}(y)=\pi_{t}(y)\cdot{\exp\big{(}\eta\mathbb{E}_{y^{\prime}\sim\pi_{t}}[u({y};y^{\prime})]\big{)}}/{Z(\pi_{t})},$ where $u(y;y^{\prime})$ is the utility function defined in Equation (1), with $y$ treated as a pure strategy.
The practical loss function of SPPO for policy update is then derived according to MWU:
@@
SPPO converges to the NE of the preference optimization game in Equation 2. However, after running multiple iterations, the deviation of the policy $\pi_{\theta}$ from $\mu$ can be large. Such deviation is particularly problematic when the preference model is only accurate at evaluating responses sampled from the reference policy (Munos et al., 2023). Furthermore, in aligning LLMs in practice, the preference model is typically a surrogate $\hat{\mathbb{P}}$, such as PairRM (Jiang et al., 2023a), which may be misspecified at some out-of-distribution responses and inaccurate due to estimation error or limited model expressiveness (PairRM is only a 0.4B model), causing over-optimization problem. Regularizing the policy optimization to a reference SFT model, which is typically trained on high-quality data (Ouyang et al., 2022), can mitigate the problem. We provide a synthetic example in Appendix D.1 to demonstrate this problem.
@@ 2.3 Regularized Preference Optimization Game with Reference Policy
To address the regularization in self-play, we adopt the objective in Nash Learning from Human Feedback (Munos et al., 2023), and extend the KL divergence regularization to a general regularization function, to penalize the deviation from the reference policy. We define a convex regularization function $R:\Delta^{\mathcal{X}}_{\mathcal{Y}}\times\Delta^{\mathcal{X}}_{\mathcal{Y}}\rightarrow(-\infty,\infty)$, where $R(\pi,\mu)$ measures the distance between $\pi$ and the reference model $\mu$, such as KL divergence $D_{\text{KL}}(\pi\|\mu)$. Denote regularization temperature as $\tau$, the objective becomes to optimize a regularized preference model by solving the NE $(\pi^{*},\pi^{*})$ of the regularized game, where the utility of max player is still $u(\pi;\pi^{\prime})=\mathbb{P}(\pi\succ\pi^{\prime})$:
@@
We provide proof of the existence and uniqueness of this NE in Appendix A.1. A few recent methods leverage Mirror Descent (MD), which is also in a self-play manner, to find a regularized NE in Equation 4 with last-iterate policy (Munos et al., 2023; Calandriello et al., 2024; Zhang et al., 2024).
However, these MD-based methods are only compatible with the reverse KL divergence regularizer, and are non-trivial to extend to general divergence. For instance, Nash-MD [^1] Throughout the paper, regularization specifically refers to the deviation of $\pi$ from $\mu$, rather than from $\pi_{t}$. addresses the reverse KL regularization of $\pi$ and $\mu$ requiring responses generated from a geometric mixture policy $\pi_{t}^{\mu}(y)\propto{\pi_{t}(y)^{1-\eta\tau}\mu(y)^{\eta\tau}}$ (Munos et al., 2023), which is inherently compatible only with reverse KL divergence:
$\pi_{t+1}=\arg\min_{\pi}-\eta\mathbb{E}_{\pi}[\nabla_{\pi}u(\pi_{t};{\color[rgb]{1,0,0}\pi^{\mu}_{t}})]+D_{\text{KL}}(\pi\|\pi_{t}^{\mu}).$ (5)
Therefore, while the LLMs optimized via existing self-play methods exhibit empirical improvement, they all have limited regularization of $\pi$ and $\mu$. The potential benefits of alternative regularization, such as adopting other $f$-divergences than reverse KL, remain unexplored.